Basic Salary of an employee is Rs. 18,000 and Allowances are Rs. 9000. According to the company’s policy casual leaves are 12 days per year and earned leaves are 24 days per year and normal working days are 22 per month.
Find the total cost of leaves as percent of gross salary.
1
Expert's answer
2019-05-21T09:34:48-0400

If we consider 22 working days in a month then there is 22*12=264 working days in a year. 12+24=36 of them are leaves. This means that a worker is effectively working 264-36=228 days a year. Now if you consider a payment of 18000 Rs. per month (22 days) you get "18000\/22 = 818.18" Rs./day when working and "9000\/22 = 409.09" Rs/day when on allowance.

As result the employer will get "409.09*36 = 14727.24" Rs. as allowances and "818.18*228= 186545.04" Rs as salary in a year.

This means that leaves cost the company


"14727.24\/(186545.04+14727.24)= 0.0732"

So 7.32%